OpenStudy (anonymous):

I need an example of an inequality using multipication and addition properties

OpenStudy (anonymous):

hmm , this is the example > x + 10 = -30 ok now lets solve it ! we add using the addition properties by adding (-10) for both sides x + 10 - 10 = -30 - 10 x = - 40 now the variable is isolated :)

OpenStudy (anonymous):

example for both multi and add. > 3x - 4 = 13 Addition Principle (+4) for both sides 3x - 4 + 4 = 13 + 4 = 3x = 17. Multiplication Principle to multiply each side by (1/3). (1/3)3x = (1/3)17 the variable is isolated. x = (17/3)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

np =D